3. Perkins Company provides the following data developed for its master budget Sales Costs: priCC $11.00 per unit Direct labor Variable overhead... Factory depreciation 3.00 per unit 4.25 per unit $0.50 per unit ation.$12,000 per month $11,000 per month Selling expensc0.25 per unit Administrative cost $9,000 per month Required: Prepare flexible budgets for sales of 20,000, 22,000 and 24,000 units. Usea contribution margin format.
Answer #1

Answer

* All working forms part of the answer

* Few things to be kept in mind while Preparing Flexible Budget:
--#1: Variable Cost changes with the change in level. Hence units are multiplied by budgeted variable cost per unit.
--#2: Fixed Cost stays the same irrespective of the amount of unit sold.

* Requirement, as asked:

Working

Contribution margin Income Statement

A

Units sold

                   20,000

                  22,000

                24,000

B = A x $ 11

Sales Revenue

$             220,000

$             242,000

$          264,000

Variable Costs:

C = A x $ 3

Direct Materials

$               60,000

$               66,000

$             72,000

D = A x $4.25

Direct Labor

$               85,000

$               93,500

$          102,000

E = A x $ 0.50

Variable Overhead

$               10,000

$               11,000

$             12,000

F = A x $ 0.25

Selling Expenses

$                  5,000

$                 5,500

$               6,000

G = C+D+E+F

Total Variable Cost

$             160,000

$             176,000

$          192,000

H = A - G

Contribution margin

$               60,000

$               66,000

$             72,000

Fixed Costs:

I

Factory Depreciation

$               12,000

$               12,000

$             12,000

J

Supervision

$               11,000

$               11,000

$             11,000

K

Administration Cost

$                  9,000

$                 9,000

$               9,000

L = I+J+K

Total Fixed Cost

$               32,000

$               32,000

$             32,000

M = H - L

Net [Operating] Income

$               28,000

$               34,000

$             40,000